# Good News for Polynomial Root-finding

### v35

@@ 8.3 The bit operation complexity of our e/i tests and root-finders
Apply Thm. 8.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^29] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates. and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 4 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle. Based on [65, Alg. 5.3] and Eqn. (8.2) obtain the following (a little larger but still reasonably small) upper bound on the bit operation cost of our subdivision root-finders,
@@
Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable). Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation B.2).
This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d-1)db$ bits and $0.25(d-1)db$ bit-operations overall for a single e/i test.
[…]
@@
Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec. B.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
@@
Bit operation cost of the shift of the variable, [132, Lemma 2.3] (cf. [65, Lemma 3.6]). Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for $\mu(s)$ of (1.8).
@@
9 Real and near real root-finding
9.1 Background, outline, and organization of our presentation
The algorithm of [67] specializes subdivision of [23] to real root-finding for univariate polynomial represented in monomial basis. Currently this is user’s choice algorithm, although we have accelerated it in [59] by narrowing the region for the search of the real roots, based on application of Newton’s polygon – by following [130] and [88, Sec.4] (cf. [89, Sec. 5]).
Unlike the customary real root-finders, next we approximate both real and near real roots, lying on and near a fixed real line segment. This covers approximation of real zeros of a polynomial given by its values or coefficients that are only available numerically with rounding errors.
Taylor’s shifts with scaling (1.4) readily reduce root-finding and root-counting for $p(x)$ in an isolated line segment to the case of the real line segment $[-1,1]$. Zhukovsky’s map of the next subsection further reduces the task to root-finding and root-counting for an auxiliary polynomial $s(z)$ on isolated unit circle $C(0,1)$ (see Remark 9.1).
If the polynomial $s(z)$ has the same root set on the unit circle $C(0,1)$ and in the annulus $A(0,\theta_{-},\theta_{+})$, for $0<\theta_{-}<1<\theta_{+}$, and if its the boundary circles are isolated, then we obtain an efficient root-counter, by applying our Alg. 3.1 to the isolated discs $D(0,\theta_{-})$ and $D(0,\theta_{+})$.
We can readily extend our subdivision algorithms to the the case of an input annulus rather than a disc. For a fixed error bound $\epsilon>0$, its complexity is dominated by the cost of the evaluation of NIR$(x)$ at $O(w\log(1/\epsilon))$ points for $w$ denoting the number of real zeros of $p(x)$ in a fixed range; realistically this number tends to be much smaller than $d$ (cf. [63, 37]).
We organize our presentation in this section as follows. In Sec. 9.2 we specify the two-way map, based on Zhukovsky’s function, from the real segment $[-1,1]$ to the unit circle $C(0,1)$ and backward. The map $[-1,1]\Longleftrightarrow C(0,1)$ can be extended to the neighborhoods of the segment and the circle (cf. Remark 9.1). We can readily compute the number $w$ of roots on the isolated unit circle $C(0,1)$ and can deflate a factor $f$ of $p$ with root set made up of the $w$ zeros of $p$ in $C(0,1)$ by extending deflation algorithms of [130] from a disc to an annulus. In Secs. 9.3 and 9.4 we extend these computations to the segment $[-1,1]$ and its neighborhood and consequently to approximation of both real and near real roots.
9.2 Zhukovsky’s function
For back and forth transition between the segment and the circle we can apply the two-to-one Zhukovsky’s function $z=J(x)$ and its one-to-two inverse, for complex variables $x$ and $z$. It maps the circle $C(0,1)$ to the segment $[-1,1]$ and vice versa:
$x=J(z):=\frac{1}{2}\Big{(}z+\frac{1}{z}\Big{)};~{}z=J^{-1}(x):=x\pm\sqrt{x^{2}-1}.$ (9.1)
9.3 A real root-finder
Algorithm 9.1.
Root-finding on a line segment.
Input:
A polynomial $p=p(x)$ of (1.1).
Output:
The number $w$ of its zeros on the unit segment $[-1,1]$ and approximations to all these zeros.
Computations:
  1. 1.
    Compute the values $v_{h}=p(\Re(\zeta_{2d}^{h}))$ of the polynomial $p$ at the Chebyshev points $\Re(\zeta_{2d}^{h})=\cos(\frac{\pi h}{2d})$ for $h=0,1,\dots,2d-1$ and $\zeta_{2d}$ of (1.10).
  2. 2.
    Interpolate to the polynomial $s(z)$ of degree $2d$ such that
    $s(z)=z^{d}p(x)~{}{\rm for}~{}x=\frac{1}{2}(z+z^{-1})$ (9.2)
    from its values
    $s(\zeta_{2d}^{h})=(-1)^{h}v_{h}~{}{\rm for}~{}h=0,1,\dots,2d-1$
    by means of Inverse DFT. [Recall that $\zeta_{2d}^{dh}=(-1)^{h}$.]
  3. 3.
    Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all roots of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$. [The zeros of $s(z)$ appear in pairs: $s(z)=0$ if $s(z)=0$ by virtue of (9.2). Furthermore, if $p(1)=0$ and/or $p(-1)=0$, then 1 and/or $-1$ are also the zeros of $g(z)$ with double multiplicity.] Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all zeros of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$. At this stage extend deflation algorithms of [130] to the the case of an input annulus about the unit circle $C(0,1)$. Output $w=0.5\deg(g)$.
  4. 4.
    By applying MPSolve, subdivision, or another root-finder approximate all $2w$ zeros of the polynomial $g(z)$. Let $y_{1},\dots,y_{w}$ denote the first $w$ of them in the order of increasing their arguments from 0.
  5. 5.
    Compute and output the values $z_{j}=\frac{1}{2}(y_{j}+\frac{1}{y_{j}})$ for $j=1,\dots,w$; these are the $w$ zeros of $p$ lying in the segment $[-1,1]$.
Correctness of this algorithm follows from (9.1) and (9.2).
Its Stage 2 is DFT. Its Stage 1 of the evaluation at Chebyshev’s points can be performed by means of the algorithm of [43] or Discrete Cosine Transform (cf. [98, Sec. 3.11 and the notes to it in Sec. 3.12]). In both cases the computation is simplified if $d$ is a power of 2; we ensure this by replacing $p$ with $x^{u}p$ for the minimal non-negative integer $u$ such that $d+u$ is a power of 2.
The complexity at stage 3 is dominated by the cost of evaluation of $p(x)$ at $O(w\log(1/\epsilon))$ points for $\epsilon$ bounding the output errors.
The converse transition from the unit circle $C(0,1)$ to the unit segment $[-1,1]$ enables us to simplify stages 4 and 5 of Alg. 9.1 by moving from $g(z)$ to a polynomial of degree $w$ whose all zeros lie in the segment $[-1,1]$. Then again we achieve this by means of evaluation and interpolation:
Algorithm 9.2.
Transition to the unit segment.
  1. 1.
    Compute the values $u_{h}=g(\zeta_{2K}^{h})$ of the polynomial $g(z)$ at the $2K$-th roots of unity $\zeta_{2K}^{h}$ for $h=0,1,\dots,K-1$ and $K>w$.
  2. 2.
    Interpolate to the polynomial $f(x)$ of degree at most $w$ from its values $f(\Re(\zeta_{2K}^{h}))=(-1)^{h}u_{h}$ at the Chebyshev points $\Re(\zeta_{2K}^{h})=\cos(\frac{\pi h}{2K})$, for $h=0,1,\dots,K-1$. [Recall that $\zeta_{2K}^{Kh}=(-1)^{h}$.]
  3. 3.
    Approximate the $w$ zeros of the polynomial $f=f(x)$ by applying MPSolve, real subdivision, e.g., those of [67], or another real polynomial root-finder.
We propose to perform steps 1 and 2 above by means of forward DFT and inverse Cosine Transforms, which we would apply to the polynomial $z^{v}g(z)$, replacing $g(x)$, where $v$ is the minimal non-negative integer such that $w+v$ is a power 2.
Remark 9.1.
Represent complex numbers as $z:=u+v\bf i$. Then Zhukovsky’s map transforms a circle $C(0,\rho)$ for $\rho\neq 1$ into the ellipse $E(0,\rho)$ whose points $(u,v)$ satisfy the following equation,
$\frac{u^{2}}{s^{2}}+\frac{v^{2}}{t^{2}}=1~{}{\rm for}~{}s=\frac{1}{2}\Big{(}\rho+\frac{1}{\rho}\Big{)},~{}~{}t=\frac{1}{2}\Big{(}\rho-\frac{1}{\rho}\Big{)}.$
Consequently it transforms the annulus $A(0,\frac{1}{\theta},\theta)$ into the domain bounded by the ellipses $E(0,\frac{1}{\theta})$ and $E(0,\theta)$, and so the circle $C(0,1)$ is $\theta$-isolated if and only if no roots of $p$ lie in the latter domain.
9.4 Real root-counting
We can simplify Alg. 9.1 and use only evaluations if we restrict our task to counting the roots that lie on or near a line segment.
Here is a high level description of this algorithm, where we do not specify the parameters $\theta$ and $q$ and assume that the input includes the polynomial $s(z)$.
Algorithm 9.3.
Root-counting on a line segment.
A real $\theta>1$ and the polynomial $s(z)$ of equation (9.2) such that the unit circle $C(0,1)$ is t $\eta$-isolated for $\eta\geq\theta^{2}$, that is, the annulus $A(0,\frac{1}{\theta^{2}},\theta^{2})$ contains no zeros of $s(z)$ except possibly some on the unit circle $C(0,1)$.
The number $w$ of the roots in the unit segment $[-1,1]$ or FAILURE.
  1. 1.
    Compute the polynomial $s_{-}(z)=s(\frac{z}{\sqrt{\theta}})$.
  2. 2.
    Choose a sufficiently large $q$ and compute Cauchy sums $\tilde{s}_{0,q}$ of the roots of $s_{-}(z)$ in the unit disc $D(0,1)$.
  3. 3.
    If the value $s_{0,q}$ is sufficiently close to an integer $j$, then output $w:=d-j$ and stop. Otherwise output FAILURE.
Let us prove correctness of the algorithm. By assumption the annulus $A(0,\frac{1}{\theta},1)$ contains no roots, except possibly some on the circle $C(0,1)$, and so $j$ is the number of the zeros of $s(z)$ in the disc $D(0,\frac{1}{\theta})$ by virtue of Observation 3.1. Clearly, the polynomial $s(z)$ of degree $2d$ has as many zeros lying in and outside the unit disc, and so precisely $2d-2j$ zeros of $s(z)$ lie on the unit circle $C(0,1)$. This is twice as many zeros as $p(x)$ has on the unit segment $[-1,1]$.
@@ 9.4 Real root-counting
Organization of Part III: We devote the next section to background. In Secs. 11 and 12 we cover our e/i tests and their extensions to $m$-tests, respectively. In Secs. 13 and 15 we estimate computational precision of our root-finders and their bit operation cost, respectively. In short Sec. 14 we extend our study of root-finding to approximation of matrix eigenvalues.
@@ 10.1 Definitions
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$, ${\rm NIR}_{t}(x){\rm NR}_{t}(x)=1$.
  • We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$.
  • We approximate NIR$(x)$ at the cost of evaluation of $p(x)$ at two points by using the first order divided differences (cf. (1.6)).
@@
  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set, or region $\mathbb{S}$ on the complex plane, respectively.
@@
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^30] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose. ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [46, Eqs. (2.2.5)–(2.2.7)])
@@
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
  • Generalize e/i tests as follows:
    Definition 10.1.
[…]
Observation 10.1.
@@ 10.2 Reverse polynomial
We will not use the following theorem but include it for the sake of completeness.
Theorem 10.1.
For $p(x)$ of (1.1) and $R>1$ it holds that
$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|,$ (10.3)
$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$ (10.4)
Proof.
To prove bound (10.3) notice that
$\frac{p(R)}{R^{d}}-p_{d}=\sum_{i=0}^{d-1}p_{i}R^{i-d}=\frac{1}{R}\sum_{i=0}^{d-1}p_{i}R^{i-d+1},$
and so
$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R}\sum_{i=0}^{d-1}|p_{i}|R^{i-d+1}\leq\max_{i<d}|p_{i}|\sum_{i=0}^{d-1}|p_{i}|R^{i-d}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|.$
Likewise, to prove bound (10.4) notice that
$Rp^{\prime}(R)-dp(R)-dR^{d-1}p_{d-1}=\sum_{i=1}^{d-1}ip_{i}R^{i-1}.$
Hence
$\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}=\frac{1}{R}\sum_{i=1}^{d-1}\frac{i}{d}p_{i}R^{i-d+1},$
and so
$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\max_{0<i<d}|p_{i}|\sum_{i=1}^{d-1}R^{i-d}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$
@@ 10.3 Root-squaring and root-lifting
10.4 The distance to a closest zero of $p$ and fast inclusion tests
Next we estimate the distance from the origin to the closest zero $z_{d}$ of $p$. The distance to its farthest zero $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
For boundons $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [77, Sec. 1.10]; in particular see Fujihara’a bounds of [39]:

### v36

@@ 12.3 The bit operation complexity of our e/i tests and root-finders
Apply Thm. 12.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^21] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates. and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 8 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
@@
for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle. Based on [68, Alg. 5.3] and Eqn. (12.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
@@
Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable). Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation 24.2).
This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d+1)db$ bits and $0.25(d+1)db$ bit-operations overall for a single e/i test.
A single subdivision iteration for Problem 1m can involve up to $4m$ e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations. This is by a factor of $m$ greater than for the root-finder of [96, 103] overall.
@@
Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec. 24.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
@@
Bit operation cost of the shift of the variable, [136, Lemma 2.3] (cf. [68, Lemma 3.6]). Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for
$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (12.5)
bounding the number of bit operations for integer multiplication modulo $2^{s}$ [53] (cf. [142, 69, 134, 135] ).
@@
Organization of Part III: We devote the next section to background. In Secs. 14 and 15 we cover our e/i tests and their extensions to $m$-tests, respectively. In Secs. 16 and 18 we estimate computational precision of our root-finders and their bit operation cost, respectively. In short Sec. 17 we extend our study of root-finding to approximation of matrix eigenvalues.
@@ 13.1 Definitions
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$. We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf. (3.2)).
@@
  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
@@
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^22] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose. ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [48, Eqs. (2.2.5)–(2.2.7)])
@@
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
  • Generalize Def. 2.1 of $\sigma$-soft e/i tests as follows: For $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^23] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. An $\ell$-test applied to the polynomial $t(y)$ of (1.3) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ. [^24] We will only use $\ell$-tests for $\ell=1$ (which are e/i tests) and $\ell=m$.
Observation 13.1.
@@ 13.3 Root-squaring and root-lifting
13.4 The distance to a closest root and fast inclusion tests
Next we estimate the distance from the origin to the closest root. The distance to the farthest root $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
For bounds $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [80, Sec. 1.10]; in particular according to Fujihara’a bounds of [41]: